The sum of three consecutive integers is one third the square of the middle number. Find the integers. Write an equation and solve.
Found 2 solutions by Alan3354, josgarithmetic:
Answer by Alan3354(69443)

The sum of three consecutive integers is one third the square of the middle number. Find the integers. Write an equation and solve.
----------------
n-1 + n + n+1 = n^2/3
n = n^2/3
3n = n^2
n = 0, n = 3
-

Answer by josgarithmetic(39617)